OpenStudy (smurfy14):

It takes ten identical pieces to form a circular track for a pair of toy racing cars. If the inside arc of each piece is 3.4 inches shorter than the outside arc, what is the width of the track?

OpenStudy (turingtest):

let r_i be the inner track and r_o be the outer track each piece has one-tenth of a circle so\[\theta=\frac{2\pi}{10}=\frac{\pi}{5}\]and for the track we have\[r_0\theta=r _i\theta+3.4\to (r_o-r_i)\theta=3.4\to r_o-r_i=\frac{3.4}{\theta}\]where r_o-r_i is the width of the track

OpenStudy (smurfy14):

thank you so much!! can you just tell me how you got 2pi over 10?

OpenStudy (turingtest):

that is the angle subtended by each piece of track the whole circle is 2pi radians a tenth of it is 2pi/10 radians (because there are ten pieces of track)

OpenStudy (smurfy14):

ohhh ok thanks

OpenStudy (turingtest):

you're welcome :)